OpenStudy (anonymous):

solve 0.6x+3≤1.4x-2 solve

OpenStudy (anonymous):

start with \[6x+30\leq 14x-20\] so make your live easier

OpenStudy (katrinakaif):

What he said.

OpenStudy (anonymous):

subtract 6x to get \[30\leq 8x-20\] add 20 to get \[50\leq 8x\] divide by 8 and be done

OpenStudy (katrinakaif):

30 + 20 <_ 14x - 6x 50 <_ 8x 50/8 <_ x 25/4 <_x

OpenStudy (saifoo.khan):

\[x \ge 6.25 \]

OpenStudy (anonymous):

saifoo has the most reliable answer and matches the equation most accurately.

OpenStudy (saifoo.khan):

hehe. thank you!! ^_^

OpenStudy (anonymous):

saifoo did you get my soybean, cormeal-protien percentage question?

OpenStudy (saifoo.khan):

That is incomplete question

OpenStudy (anonymous):

It went like this-hold on

OpenStudy (katrinakaif):

Yes..It seems to have no useful information to solve the problem,

OpenStudy (saifoo.khan):

Right.

OpenStudy (saifoo.khan):

that stuff is useless.. hehe

OpenStudy (anonymous):

Soybean meal is 16% protien, cornmeal is 8% protien.How many pounds of each shoud be mixed together to get a 320 pound mixture that is 12% protien?

OpenStudy (katrinakaif):

S: number of pounds of soybean C: number of pounds of cornmeal .16(s) + .8(c) = 320(.12) s + c = 320

OpenStudy (anonymous):

so that would be exactly how many pounds of each?

OpenStudy (saifoo.khan):

Answer: 160 lbs of soybeans, 160 lbs of cornmeal.

OpenStudy (katrinakaif):

Well turn the equation to something we can work with no messy decimals, so multiply the top equation by 100 resulting: 16s + 8c = 320 (12) ---> 3840 Now subsitute the 's + c = 320' equation for one of the variables, s = 320 - c Subsitute: 16(320 - c) + 8c = 3840 Solve for c, then subsitute that value as 'c' in the equation s + c = 320

OpenStudy (anonymous):

which equals how many pounds of each?

OpenStudy (anonymous):

lol

OpenStudy (katrinakaif):

Let me check saif's answer.

OpenStudy (anonymous):

your killing me

OpenStudy (katrinakaif):

Come on man. Its easy multiplication..

OpenStudy (anonymous):

as easy as how many pounds each and checking saifs answer, saif, where the hell are you?

OpenStudy (katrinakaif):

You know what. You do the work you lazy lump.

OpenStudy (anonymous):

looks like you couldn't answer it eeither bluffer

OpenStudy (katrinakaif):

Looks like you forgot how to do multiplication and division. Even Satellite has a decent explanation to lead you to the answer.

OpenStudy (anonymous):

if you wish to do no work whatsoever, just type the dang thing in and read off the answer http://www.wolframalpha.com/input/?i=0.16x%2B.08%28320-x%29%3D.12*320